OpenStudy (anonymous):

A standard deck of playing cards contains 52 cards, equally divided among four suits (hearts, diamonds, clubs, and spades). Each suit has the cards 2 through 10, as well as a jack, a queen, a king, and an ace. If the 3 of spades is drawn from a standard deck and is not replaced, what is the probability that the next card drawn is a spade OR a king? A. 1/17 B. 16/51 C. 4/17 D. 5/17

OpenStudy (anonymous):

Let's see... There would be 51 cards left, including 12 spades and 4 kings. One of the kings is a spade. 12 + 4 = 16, but since one of the kings is a spade, 16 - 1 = 15. So the probability of drawing a spade OR king next is 15/51

OpenStudy (accessdenied):

A standard deck has 13 of each suit. Hearts: 2 3 4 5 6 7 8 9 10 J Q K A Diamonds: 2 3 4 5 6 7 8 9 10 J Q K A Clubs: 2 3 4 5 6 7 8 9 10 J Q K A Spades: 2 3 4 5 6 7 8 9 10 J Q K A We removed the 3 of spades. Our deck of cards now has 51 cards, including only 12 spades. Here I mark the outcomes we want with brackets: Hearts: 2 3 4 5 6 7 8 9 10 J Q [K] A Diamonds: 2 3 4 5 6 7 8 9 10 J Q [K] A Clubs: 2 3 4 5 6 7 8 9 10 J Q [K] A Spades: [2 4 5 6 7 8 9 10 J Q] K [A] "spades or king" => cannot be king and spades We are looking at two events that are not mutually exclusive (we can draw a card that matches both conditions, which would not fit the "or") There is a 12/51 chance of drawing a spades and a 4/51 chance of drawing a king. We cannot include the fourth king of spades as an option (1/51). 4/51 + 12/51 - 1/51 = 15/51 You'd just have to simplify 15/51